Astronomers have discovered a planetary system orbiting a star, which is at a distance of 5.2 × 1018 m from the earth. One planet is believed to be located at a distance of 1.9 × 1010 m from the star. Using visible light with a vacuum wavelength of 557 nm, what is the minimum necessary aperture diameter that a telescope must have so that it can resolve the planet and the star?
Answer #1

The angular distance subtended by the planet orbit as seen from Earth is

\theta\sim\frac{r}{d}=\frac{1.9\cdot 10^{10}}{5.2\cdot 10^{18}}\sim4\cdot 10^{-9}rad

The telescope must have an aperture so that its PSF due to diffraction is smaller than this value, thus

\frac{\lambda}{A}<\theta

thus

A>\frac{\lambda}{\theta}\sim139m